Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Geoffrey Stewart, Ying Tang
Ongoing Systems Updates
- Geoffrey met his new neighbor – a Yankee living out in Newberry!
- Culligan water saga – no water in Systems at the moment. If we don’t hear from them by 3PM today, we will try, try again.
- Orders have been going out.
- Requests for Web authoring software – we’ve added a disclaimer to SysRequest that these are not going to be purchased until the Library knows what it’s doing.
- Items ordered on PO – unsure how those are proceeding.
- Updates – close to complete. IE patch issued on Friday has been pushed. Restart tomorrow, check update site to verify patch has been installed.
- Will is rebuilding the patch server box.
- Major spyware infestation in Resource Services. Took a very long time to fix. Basic result is a fountain of pop-up ads, rendering the machine unusable. Bill is testing Spybot at home. It will be used on staff machines only, if we decide to use it.
- Use of IE? Firefox work just as well. Bill is still leaning towards instructing the liaisons to stop the use of IE and install Firefox as the default browser on all staff machines. We’ll see how the IE patch goes, but it really is not a risk worth taking. Firefox is faster and cleaner. We will distribute it via the server.
Software Projects Status
- Two major disruptions to campus security. Internally, a workstation in Education was compromised, resulting in the installation of a FTP server on a staff machine. It was used to get into a German server. We took it off the network immediately. Will is going to check it out after the meeting today. The local admin password in Education had to be changed.
- CLAS listserv got whacked. It was stopped here, but got through at CLAS. May be getting by McAfee.
- ALEPH printing – status quo.
- Public workstations – Geoffrey is cleaning up the second draft of the documentation. IE or Firefox? Public will remain IE for now. Geoffrey is working on some small details in getting it to work. Bill wants it ready to go for the Liaison Meeting on Thursday.
- POS/Inventory – met with Rob & Russ, they liked all of the changes. They requested one extra report: it’s done. They had on their wish list a request for a Web page so that staff can see what they have in the inventory.
- Spine Label – one change – dymo printer designates any printer as the strip printer. Had to change location of the temporary file.
- Ying and Mark have come up with a nice database design for DLC Tracking. Ying had to add new data and finish C-Sharp and STOR procedures.
Student and Departmental Happenings
- Per Tom, the new Library homepage was supposed to be in place today. However, it was decided that we would wait until finals are over. There is a link to the conversion on the current homepage, so check it out and leave feedback. New page should be up by next week.
- Departmental entry pages (at the very least) should be converted by now. The rest of the conversion is expected to continue through the fall semester.
- Tape copying is well underway.
- Students – this is the last week of Summer B. Adrianne will be leaving, but has referred her roommate for hiring. Grace will stay and Lisa Gaul should be returning for the fall.
Back to the Systems Homepage
Last updated August 10, 2004
by Debra Fetzer